# literature.bib

### v4

@@ 1. Introduction
in the new variational framework of [AV22variational] by Agresti and the second author. This framework, the critical variational setting, has been developed to extend the classical variational approach to stochastic evolution equations originating from [bensoussan72], [pardoux], [krylov81]. In the classical variational approach, the drift and diffusion coefficients $A$ and $B$ need to satisfy several conditions to ensure well-posedness of (1.1). The usual weak monotonicity condition is especially restrictive. It is therefore no surprise that efforts have been made to weaken the monotonicity condition, e.g. in [liurockner15, §5.2] and very recently [rockner22] with a much weaker local monotonicity condition. One of the advantages of the critical variational setting of [AV22variational] used in this paper, is that no form of monotonicity is assumed. In return, $A$ and $B$ are of a slightly less (but still very) general form:
@@
for $t\in{\mathbb{R}}_{+}$ and $v\in V$, where $(V,H,V^{*})$ is a Gelfand triple belonging to the stochastic evolution equation. That is, $(A,B)$ contains a quasilinear part $(A_{0},B_{0})$ and a semilinear part $(F,G)$ and it is assumed that both parts satisfy certain critical local Lipschitz conditions, where the Lipschitz constant may depend arbitrarily on $\|v\|_{H}$ and, allowing even more flexibility, polynomially on interpolation norms $\|v\|_{V_{\beta}}$, where $V_{\beta}=[V^{*},V]_{\beta}$ denotes the complex interpolation space. Besides the absence of any monotonicity assumption, another major improvement of the critical variational setting is the weakening of the usual growth conditions on the diffusion coefficient $B$, allowing e.g. for gradient noise. Lastly, a special feature is that critical nonlinearities are allowed (see (2.6)), which is not the case in other settings. The critical variational setting covers many semilinear and some quasilinear equations that were not covered by more classical variational settings. In particular, this holds for many equations that require an (analytically) strong setting, in which monotonicity often fails, for example the Cahn–Hilliard equation, the tamed Navier–Stokes equations and the Allen–Cahn equation. See [AV22variational, §5] for details. The exact assumptions in the critical variational setting can be found in Section 2. Finally, it should be stressed that unlike the settings in [rockner22], the critical variational setting does not require a compact (Sobolev) embedding $V\hookrightarrow H$ and is thus suited to treat equations on unbounded spatial domains.
@@
The goal of this paper is to establish the LDP for solutions to small-noise stochastic evolution equations in the critical variational setting. Large deviations have been studied for SPDEs in many different frameworks. The first results for SPDEs were inspired by the pioneering paper for SDEs by Freidlin and Wentzell [freidlin12] (see also [dembo10, §5.6]), relying on discretizations and the contraction principle. These techniques were extended to several SPDE settings with Gaussian noise, notably in [daprato14], [peszat94], [cerrai04] (stochastic reaction-diffusion equations), [chow92] (semilinear parabolic equations) and [rockner06] (stochastic porous media equations). However, for less regular $A$ and $B$ such techniques are difficult to use in general settings. In 2001, Budhiraja and Dupuis proved a substantially generalized contraction principle, the so-called weak convergence approach to large deviations [budhidupuis01]. This approach turned out to be extremely powerful for SPDEs and subsequently, it was applied to many SPDEs with less regular coefficients, e.g. in [sritharansundar06], [chueshovmillet10] (2D Navier–Stokes and hydrodynamical models), [duanmillet09] (Boussinesq equations), as well as [renzhang08] and [liu09] (general classical variational settings). A more detailed discussion on applications to fluid dynamics can be found below.
@@
Also for the recent variational settings with even weaker conditions on the coefficients $A$ and $B$, the weak convergence approach has led to new LDP proofs. In [hongliliu21] the LDP is obtained for McKean-Vlasov quasilinear stochastic evolution equations, in [pan22] for a setting from [rockner22], in [kumarmohan22] for the same setting extended to Lévy noise, and most recently, [pan24] obtained the LDP for the strongest setting of [rockner22]. The latter allows flexible growth bounds on $B$, including gradient noise. Still, the combination of flexible growth of $B$ and unbounded spatial domains (that is, no compact embedding $V\hookrightarrow H$) has not been covered in any of the papers so far. The main improvement of our work is that we allow for both. In fact, no additional bounds on $A$ and $B$ are assumed for the LDP apart from those in [AV22variational] required for well-posedness, nor do we assume a compact embedding in the Gelfand triple. New techniques are used to replace the usual compactness arguments. The paper contains new approaches for
@@
    Navier–Stokes equations for $d=2$ [AV24SNS, App. A],
    tamed Navier–Stokes equations for $d=3$ [AV22variational, §5.2],
    Cahn–Hilliard equation for $d=1,2$ [AV22variational, §5.1],
    Swift–Hohenberg equations for $d=1,2,3$ [AV22variational, §5.6],
@@
    • Allen–Cahn equations [AV22variational, §5.4],
    • symbiotic Lotka–Volterra equations [AV24reacdiff-arXiv, Th. 3.11],
    • coagulation equations [AV24reacdiff-arXiv, Th. 3.9].
@@
The LDP for the 2D Navier–Stokes equations with gradient noise and unbounded domains was already considered in the pioneering work [sritharansundar06]. However, the proof of [sritharansundar06, Lem. 4.5] is incomplete – specifically, the argument in the final line. The gap was already indicated in [duanmillet09, p. 2054] and concerns the compactness of the sublevel sets of the rate function. In [chueshovmillet10, duanmillet09], the gap is resolved for the 2D Navier–Stokes and Boussinesq equations under the assumption that the noise is gradient-free (see [chueshovmillet10, Th. 3.2], [duanmillet09, Ass. A Bis, p. 2072]), but the gradient noise case has remained open since then. The possibly unbounded spatial domains rule out alternative arguments based on compact Sobolev embeddings. We have now covered the gradient noise case in Theorems 5.2 and 5.4 – extending the results of [chueshovmillet10, duanmillet09] and completing the proof of [sritharansundar06].
Another application that we would like to highlight are the 3D tamed Navier–Stokes equations, for which a large deviation principle was established in [RoZhZh]. However, gradient noise was not considered in that work, and it is far from straightforward to extend their approach to settings where such noise is present. Our main result, Theorem 2.6, now includes the gradient noise case and also applies to a broad class of other models (see the list above).
@@ 2.1. The critical variational setting
[…]
@@
For convenience of the reader, we recall that in applications, one does not work with the abstract dual $V^{*}$ but with a space $V^{\prime}$ which, under some assumptions and with the correct duality pairing, is isomorphic to $V^{*}$, see also [krylov81, p. 1244]. One starts with reflexive Banach (or Hilbert) (sub)spaces $V\subset H\subset V^{\prime}$, where each inclusion is dense and continuous and one defines $j\colon H\hookrightarrow V^{*}\colon x\mapsto(x,\cdot)_{H}$. Then, provided that
@@
The equivalence follows from [rynne08, Th. 4.48], density of $H\subset V^{\prime}$ and continuity of $j_{1}$, and density of $\mathrm{Im}(j_{1})\subset V^{*}$. The latter holds since $j_{1}(V^{\prime})\supset j_{1}(H)=j(H)$ and one can verify that $j(H)$ is dense in $V^{*}$ using reflexivity of $V$. In conclusion, provided that $V\subset H\subset V^{\prime}$ continuously and densely, one only has to verify (2.1) and (2.2) to have $j_{1}\colon V^{\prime}\cong V^{*}$.
@@
[…]
See also [AV22variational, Ex. 2.1, Ex. 2.2].
Recall that $H=[V^{*},V]_{\frac{1}{2}}$ [arendt02, §5.5.2] and the following interpolation estimate holds for $\beta\in(\frac{1}{2},1)$:
@@
Since strong solutions are required to be strongly measurable, see Definition 2.1 below, one can assume without loss of generality that $V$ and $H$ are separable, see also [krylov81, p. 1244]. Thus, from now on we assume that $V$ and $H$ are separable.
As mentioned in the introduction, we work with the critical variational setting from [AV22variational]. We consider stochastic evolution equations of the form
@@
If $\Phi:[0,T]\times\Omega\to\mathcal{L}_{2}(U,H)$ is strongly progressively measurable and $\Phi\in L^{2}(0,T;\mathcal{L}_{2}(U,H))$ a.s., then one can define the stochastic integral $\int_{0}^{t}\Phi(s)\,\mathrm{d}W(s)$ for $t\in[0,T]$, see [NVW15, §5.4 $(p=0)$].
@@
For the weak convergence approach to large deviations it is necessary to let $A$ and $B$ be defined on +×${\mathbb{R}}_{+}\times V$ rather than +×${\mathbb{R}}_{+}\times\Omega\times V$, meaning that stochasticity enters $A$ and $B$ through the solution $u$ in (2.4) and not separately. Also, the initial value $x$ in (1.1) has to be deterministic. Other than that, we make exactly the same assumptions as those required for global well-posedness [AV22variational, Th. 3.5]. Let us introduce these assumptions.
@@
Because the coefficients are defined on +×${\mathbb{R}}_{+}\times V$ instead of +×${\mathbb{R}}_{+}\times\Omega\times V$, the measurability in Assumption 2.2(1) is different than in [AV22variational, Assumption 3.1]. However, $(A,B)$ satisfies our assumption if and only if $(\bar{A},\bar{B})$ satisfies [AV22variational, Assumption 3.1], where $\bar{A}(t,\omega,v)\coloneqq A(t,v)$ and $\bar{B}(t,\omega,v)\coloneqq B(t,v)$ are trivial extensions.
The operators $A_{0}$ and $B_{0}$ are of leading (differential) order and of quasilinear form. In the semilinear case, they are linear in the sense that $A_{0}(t,u)$ and $B_{0}(t,u)$ do not depend on $u$.
@@
From (2.3) and Assumption 2.2(3), it is clear that $\|F(t,v)\|_{V^{*}}+|\!|\!|G(t,v)|\!|\!|_{H}\leq\tilde{C}_{\|v\|_{H},T}(1+\|v\|_{V})$ for all $t\in[0,T]$ if $\|v\|_{H}\leq n$, where $\tilde{C}_{\|v\|_{H},T}$ is a constant. Thus we have integrability of $F(\cdot,u(\cdot))$ and $G(\cdot,u(\cdot))$ if $u\in\mathrm{MR}(0,T)$.
In [AV22variational, Th. 3.3] it is shown that under Assumption 2.2, there exists a unique local solution to (2.4). In [AV22variational, Th. 3.5], this is extended to a global well-posedness result under a coercivity condition on $(A,B)$. The next result follows from [AV22variational, Th. 3.5].
@@
Energy estimates can also be found in [AV22variational, Th. 3.5], but these will not be used. More general theory in an $L^{p}$-setting was developed in [AV22nonlinear1] and [AV22nonlinear2].
@@ 2.2. Statement of the main result
    $I$ has compact sublevel sets,
@@
To have $u^{\psi}$ appearing in (2.9) well-defined, (2.8) needs to be (globally) well-posed. In Section 3, we prove that this is the case. Finally, we recall that the LDP is equivalent to the Laplace principle [dupuis97, Def. 1.2.2, Th. 1.2.1, Th. 1.2.3]. The weak convergence approach from [budhidupuis01] offers sufficient conditions for the latter, hence for the LDP. The approach is stated in Subsection 4.1, after which we apply it to prove Theorem 2.6 in the remainder of Section 4.
@@ 3. Well-posedness of the skeleton equation
Unfortunately, well-posedness cannot be proved at once. Instead, we first achieve well-posedness of an appropriate linearized version of the skeleton equation in Corollary 3.5, together with a maximal regularity estimate. Then, we can borrow the strategies from [HNVWvolume3, Chap. 18], [pruss17addendum], [pruss18JDE], [pruss16]. That is, we use the maximal regularity estimate of Corollary 3.5 for the linearized equation in a fixed point argument, yielding existence of a local solution to the skeleton equation in Theorem 3.7. Finally, we extend to a global solution in Theorem 3.16, making use of a blow-up criterion. Uniqueness will be obtained along the way.
@@ 3.1. Linearized skeleton equation
where $A_{0}$ and $B_{0}$ are as in Assumption 2.2 and $\bar{f}\in L^{2}(0,T;V^{*})$, $\bar{g}\in L^{2}(0,T;\mathcal{L}_{2}(U,H))$. In this subsection we prove well-posedness of (3.1) using the method of continuity [HNVWvolume3, Lem. 16.2.2], together with a suitable maximal regularity estimate. We prove it for more general equations as this does not require any more effort and makes the exposition more transparent. Let us introduce spaces $S$ and $E$ that will be used in the method of continuity.
@@
Note that $S$ is a Banach space [HNVWvolume1, Prop. C.1.3] and $S\hookrightarrow L^{1}(0,T;V^{*})$. Moreover, we define
$E\coloneqq\{u\in\mathrm{MR}(0,T):u\text{ is weakly differentiable, }u^{\prime}\in S\},\quad\|u\|_{E}\coloneqq\|u\|_{\mathrm{MR}(0,T)}+\|u^{\prime}\|_{S}.$
@@
Dealing with the sum space $S$ is not standard in part of the literature. However, it is covered excellently in Pardoux’ thesis [pardoux].
The following proposition is a direct consequence of [pardoux, Th. 2.1].
@@
Let $\bar{A}\colon[0,T]\to\mathcal{L}(V,V^{*})$ be such that $[0,T]\to V^{*}\colon t\mapsto\bar{A}(t)v$ is strongly Borel measurable for all $v\in V$ and suppose that $a_{T}\coloneqq\sup_{t\in[0,T]}\|\bar{A}(t)\|_{\mathcal{L}(V,V^{*})}<\infty$. Suppose that there exists $\theta>0$ such that for all $t\in[0,T]$ and $v\in V$:
@@
Write $h=f+g$ with $f\in L^{2}(0,T;V^{*})$, $g\in L^{1}(0,T;H)$. We apply [pardoux, Th. 2.2]. Since $u\in\mathrm{MR}(0,T)$ is a strong solution, we have $u(t)=x+\int_{0}^{t}v(s)\,\mathrm{d}s$ with $v\coloneqq h(\cdot)-\bar{A}(\cdot)u(\cdot)\in S=L^{2}(0,T;V^{*})+L^{1}(0,T;H)$. Hence, the chain rule (A.2) and (3.3) yield for all $t\in[0,T]$:
@@
We use the method of continuity [HNVWvolume3, Lem. 16.2.2]. Define $A_{0}\in\mathcal{L}(V,V^{*})$ by $A_{0}v\coloneqq\theta(\cdot,v)_{V}$. For $\lambda\in[0,1]$, put
@@
Clearly, $L_{\lambda}$ is linear. We show that $L_{\lambda}\in\mathcal{L}(E,S\times H)$ and that $[0,1]\to\mathcal{L}(E,S\times H)\colon\lambda\mapsto L_{\lambda}$ is continuous.
@@
$\|A_{0}(\cdot,w(\cdot))u(\cdot)\|_{L^{2}(0,T;V^{*})}\leq C_{n,T}(1+n)\|u\|_{L^{2}(0,T;V)}<\infty$
@@
where $\alpha\coloneqq C_{n,T}(1+n)(1+\|\psi\|_{L^{2}(0,T;U))})$.
@@
As $u|_{[0,\tilde{T}]}$ is a strong solution to (3.1) on $[0,\tilde{T}]$, (3.4) and (3.9) yield

### v5

# Large Deviations for Stochastic Evolution Equations in the Critical Variational Setting

@@ 1. Introduction
in the new variational framework of [undefe] by Agresti and the second author. This framework, the critical variational setting, has been developed to extend the classical variational approach to stochastic evolution equations originating from [undefg], [undefag], [undefy]. In the classical variational approach, the drift and diffusion coefficients $A$ and $B$ need to satisfy several conditions to ensure well-posedness of (1.1). The usual weak monotonicity condition is especially restrictive. It is therefore no surprise that efforts have been made to weaken the monotonicity condition, e.g. in [undefab, §5.2] and very recently [undefan] with a much weaker local monotonicity condition. One of the advantages of the critical variational setting of [undefe] used in this paper, is that no form of monotonicity is assumed. In return, $A$ and $B$ are of a slightly less (but still very) general form:
@@
for $t\in{\mathbb{R}}_{+}$ and $v\in V$, where $(V,H,V^{*})$ is a Gelfand triple belonging to the stochastic evolution equation. That is, $(A,B)$ contains a quasilinear part $(A_{0},B_{0})$ and a semilinear part $(F,G)$ and it is assumed that both parts satisfy certain critical local Lipschitz conditions, where the Lipschitz constant may depend arbitrarily on $\|v\|_{H}$ and, allowing even more flexibility, polynomially on interpolation norms $\|v\|_{V_{\beta}}$, where $V_{\beta}=[V^{*},V]_{\beta}$ denotes the complex interpolation space. Besides the absence of any monotonicity assumption, another major improvement of the critical variational setting is the weakening of the usual growth conditions on the diffusion coefficient $B$, allowing e.g. for gradient noise. Lastly, a special feature is that critical nonlinearities are allowed (see (2.6)), which is not the case in other settings. The critical variational setting covers many semilinear and some quasilinear equations that were not covered by more classical variational settings. In particular, this holds for many equations that require an (analytically) strong setting, in which monotonicity often fails, for example the Cahn–Hilliard equation, the tamed Navier–Stokes equations and the Allen–Cahn equation. See [undefe, §5] for details. The exact assumptions in the critical variational setting can be found in Section 2. Finally, it should be stressed that unlike the settings in [undefan], the critical variational setting does not require a compact (Sobolev) embedding $V\hookrightarrow H$ and is thus suited to treat equations on unbounded spatial domains.
@@
The goal of this paper is to establish the LDP for solutions to small-noise stochastic evolution equations in the critical variational setting. Large deviations have been studied for SPDEs in many different frameworks. The first results for SPDEs were inspired by the pioneering paper for SDEs by Freidlin and Wentzell [undefq] (see also [undefm, §5.6]), relying on discretizations and the contraction principle. These techniques were extended to several SPDE settings with Gaussian noise, notably in [undefl], [undefah], [undefi] (stochastic reaction-diffusion equations), [undefj] (semilinear parabolic equations) and [undefao] (stochastic porous media equations). However, for less regular $A$ and $B$ such techniques are difficult to use in general settings. In 2001, Budhiraja and Dupuis proved a substantially generalized contraction principle, the so-called weak convergence approach to large deviations [undefh]. This approach turned out to be extremely powerful for SPDEs and subsequently, it was applied to many SPDEs with less regular coefficients, e.g. in [undefar], [undefk] (2D Navier–Stokes and hydrodynamical models), [undefn] (Boussinesq equations), as well as [undefal] and [undefaa] (general classical variational settings). A more detailed discussion on applications to fluid dynamics can be found below.
@@
Also for the recent variational settings with even weaker conditions on the coefficients $A$ and $B$, the weak convergence approach has led to new LDP proofs. In [undefs] the LDP is obtained for McKean-Vlasov quasilinear stochastic evolution equations, in [undefae] for a setting from [undefan], in [undefz] for the same setting extended to Lévy noise, and most recently, [undefaf] obtained the LDP for the strongest setting of [undefan]. The latter allows flexible growth bounds on $B$, including gradient noise. Still, the combination of flexible growth of $B$ and unbounded spatial domains (that is, no compact embedding $V\hookrightarrow H$) has not been covered in any of the papers so far. The main improvement of our work is that we allow for both. In fact, no additional bounds on $A$ and $B$ are assumed for the LDP apart from those in [undefe] required for well-posedness, nor do we assume a compact embedding in the Gelfand triple. New techniques are used to replace the usual compactness arguments. The paper contains new approaches for
@@
    Navier–Stokes equations for $d=2$ [undefd, App. A],
    tamed Navier–Stokes equations for $d=3$ [undefe, §5.2],
    Cahn–Hilliard equation for $d=1,2$ [undefe, §5.1],
    Swift–Hohenberg equations for $d=1,2,3$ [undefe, §5.6],
@@
    • Allen–Cahn equations [undefe, §5.4],
    • symbiotic Lotka–Volterra equations [undefc, Th. 3.11],
    • coagulation equations [undefc, Th. 3.9].
@@
The LDP for the 2D Navier–Stokes equations with gradient noise and unbounded domains was already considered in the pioneering work [undefar]. However, the proof of [undefar, Lem. 4.5] is incomplete – specifically, the argument in the final line. The gap was already indicated in [undefn, p. 2054] and concerns the compactness of the sublevel sets of the rate function. In [undefk, undefn], the gap is resolved for the 2D Navier–Stokes and Boussinesq equations under the assumption that the noise is gradient-free (see [undefk, Th. 3.2], [undefn, Ass. A Bis, p. 2072]), but the gradient noise case has remained open since then. The possibly unbounded spatial domains rule out alternative arguments based on compact Sobolev embeddings. We have now covered the gradient noise case in Theorems 5.2 and 5.4 – extending the results of [undefk, undefn] and completing the proof of [undefar].
Another application that we would like to highlight are the 3D tamed Navier–Stokes equations, for which a large deviation principle was established in [undefap]. However, gradient noise was not considered in that work, and it is far from straightforward to extend their approach to settings where such noise is present. Our main result, Theorem 2.6, now includes the gradient noise case and also applies to a broad class of other models (see the list above).
@@ 2.1. The critical variational setting
We let $(V,H,V^{*})$ be a Gelfand triple of real Hilbert spaces. That is, $(V,(\cdot,\cdot)_{V})$ and $(H,(\cdot,\cdot)_{H})$ are real Hilbert spaces such that there exists a continuous and dense embedding $\iota\colon V\hookrightarrow H$. Then, $j\colon H\hookrightarrow V^{*}\colon x\mapsto(x,\iota(\cdot))_{H}$ is a continuous embedding and $j(H)$ is dense in $V^{*}$ by reflexivity of $V$ ($j=\iota^{*}$ under Riesz’ identification $H\cong H^{*}$). From now on we identify $x\in V$ with $\iota(x)\in H$ and $x\in H$ with $j(x)\in V^{*}$. Then, if $\langle\cdot,\cdot\rangle$ denotes the duality pairing between the abstract dual $V^{*}$ and $V$, one has
@@
For convenience of the reader, we recall that in applications, one does not work with the abstract dual $V^{*}$ but with a space $V^{\prime}$ which, under some assumptions and with the correct duality pairing, is isomorphic to $V^{*}$, see also [undefy, p. 1244]. One starts with reflexive Banach (or Hilbert) (sub)spaces $V\subset H\subset V^{\prime}$, where each inclusion is dense and continuous and one defines $j\colon H\hookrightarrow V^{*}\colon x\mapsto(x,\cdot)_{H}$. Then, provided that
@@
The equivalence follows from [undefaq, Th. 4.48], density of $H\subset V^{\prime}$ and continuity of $j_{1}$, and density of $\mathrm{Im}(j_{1})\subset V^{*}$. The latter holds since $j_{1}(V^{\prime})\supset j_{1}(H)=j(H)$ and one can verify that $j(H)$ is dense in $V^{*}$ using reflexivity of $V$. In conclusion, provided that $V\subset H\subset V^{\prime}$ continuously and densely, one only has to verify (2.1) and (2.2) to have $j_{1}\colon V^{\prime}\cong V^{*}$.
@@
$V=H^{1}({\mathbb{R}}^{d}),\;H=L^{2}({\mathbb{R}}^{d}),\;V^{\prime}=H^{-1}({\mathbb{R}}^{d})=V^{*}$ $\text{(weak setting)},$
$V=H^{2}({\mathbb{R}}^{d}),\;H=H^{1}({\mathbb{R}}^{d}),\;V^{\prime}=L^{2}({\mathbb{R}}^{d})\cong V^{*}$ $\text{(strong setting)}.$
See also [undefe, Ex. 2.1, Ex. 2.2].
Recall that $H=[V^{*},V]_{\frac{1}{2}}$ [undeff, §5.5.2] and the following interpolation estimate holds for $\beta\in(\frac{1}{2},1)$:
@@
Since strong solutions are required to be strongly measurable, see Definition 2.1 below, one can assume without loss of generality that $V$ and $H$ are separable, see also [undefy, p. 1244]. Thus, from now on we assume that $V$ and $H$ are separable.
As mentioned in the introduction, we work with the critical variational setting from [undefe]. We consider stochastic evolution equations of the form
@@
If $\Phi:[0,T]\times\Omega\to\mathcal{L}_{2}(U,H)$ is strongly progressively measurable and $\Phi\in L^{2}(0,T;\mathcal{L}_{2}(U,H))$ a.s., then one can define the stochastic integral $\int_{0}^{t}\Phi(s)\,\mathrm{d}W(s)$ for $t\in[0,T]$, see [undefad, §5.4 $(p=0)$].
@@
For the weak convergence approach to large deviations it is necessary to let $A$ and $B$ be defined on ${\mathbb{R}}_{+}\times V$ rather than ${\mathbb{R}}_{+}\times\Omega\times V$, meaning that stochasticity enters $A$ and $B$ through the solution $u$ in (2.4) and not separately. Also, the initial value $x$ in (1.1) has to be deterministic. Other than that, we make exactly the same assumptions as those required for global well-posedness [undefe, Th. 3.5]. Let us introduce these assumptions.
@@
Because the coefficients are defined on ${\mathbb{R}}_{+}\times V$ instead of ${\mathbb{R}}_{+}\times\Omega\times V$, the measurability in Assumption 2.2(1) is different than in [undefe, Ass. 3.1]. However, $(A,B)$ satisfies our assumption if and only if $(\bar{A},\bar{B})$ satisfies [undefe, Ass. 3.1], where $\bar{A}(t,\omega,v)\coloneqq A(t,v)$ and $\bar{B}(t,\omega,v)\coloneqq B(t,v)$ are trivial extensions.
$A_{0}$ and $B_{0}$ determine the leading differential order of the drift and the noise, respectively, and constitute the principal part of the quasilinear equation. In the semilinear case, $A_{0}(t,u)$ and $B_{0}(t,u)$ do not depend on $u$, rendering the principal part linear.
@@
[…]
In [undefe, Th. 3.3] it is shown that under Assumption 2.2, there exists a unique local solution to (2.4). In [undefe, Th. 3.5], this is extended to a global well-posedness result under a coercivity condition on $(A,B)$. The next result follows from [undefe, Th. 3.5].
@@
Energy estimates can also be found in [undefe, Th. 3.5], but these will not be used. More general theory in an $L^{p}$-setting was developed in [undef, undefa].
@@ 2.2. Statement of the main result
    II has compact sublevel sets,
@@
To have $u^{\psi}$ appearing in (2.9) well-defined, (2.8) needs to be (globally) well-posed. In Section 3, we prove that this is the case. Finally, we recall that the LDP is equivalent to the Laplace principle [undefo, Def. 1.2.2, Th. 1.2.1, Th. 1.2.3]. The weak convergence approach from [undefh] offers sufficient conditions for the latter, hence for the LDP. The approach is stated in Subsection 4.1, after which we apply it to prove Theorem 2.6 in the remainder of Section 4.
@@ 3. Well-posedness of the skeleton equation
Unfortunately, well-posedness cannot be proved at once. Instead, we first achieve well-posedness of an appropriate linearized version of the skeleton equation in Corollary 3.5, together with a maximal regularity estimate. Then, we can borrow the strategies from [undefu, Chap. 18], [undefak], [undefaj], [undefai]. That is, we use the maximal regularity estimate of Corollary 3.5 for the linearized equation in a fixed point argument, yielding existence of a local solution to the skeleton equation in Theorem 3.7. Finally, we extend to a global solution in Theorem 3.16, making use of a blow-up criterion. Uniqueness will be obtained along the way.
@@ 3.1. Linearized skeleton equation
where $A_{0}$ and $B_{0}$ are as in Assumption 2.2 and $\bar{f}\in L^{2}(0,T;V^{*})$, $\bar{g}\in L^{2}(0,T;\mathcal{L}_{2}(U,H))$. In this subsection we prove well-posedness of (3.1) using the method of continuity [undefu, Lem. 16.2.2], together with a suitable maximal regularity estimate. We prove it for more general equations as this does not require any more effort and makes the exposition more transparent. Let us introduce spaces $S$ and $E$ that will be used in the method of continuity.
@@
Note that $S$ is a Banach space [undeft, Prop. C.1.3] and $S\hookrightarrow L^{1}(0,T;V^{*})$. Moreover, we define
[…]
@@
Dealing with the sum space $S$ is not standard in part of the literature. However, it is covered excellently in Pardoux’ thesis [undefag].
The following proposition is a direct consequence of [undefag, Th. 2.1].
@@
Let $\bar{A}\colon[0,T]\to\mathcal{L}(V,V^{*})$ be such that $0,T]\to V^{*}\colon t\mapsto\bar{A}(t)v$ is strongly Borel measurable for all $v\in V$ and suppose that $a_{T}\coloneqq\sup_{t\in[0,T]}\|\bar{A}(t)\|_{\mathcal{L}(V,V^{*})}<\infty$. Suppose that there exists $\theta>0$ such that for all $t\in[0,T]$ and $v\in V$:
@@
Write $h=f+g$ with $f\in L^{2}(0,T;V^{*})$, $g\in L^{1}(0,T;H)$. We apply [undefag, Th. 2.2]. Since $u\in\mathrm{MR}(0,T)$ is a strong solution, we have $u(t)=x+\int_{0}^{t}v(s)\,\mathrm{d}s$ with $v\coloneqq h(\cdot)-\bar{A}(\cdot)u(\cdot)\in S=L^{2}(0,T;V^{*})+L^{1}(0,T;H)$. Hence, the chain rule (A.2) and (3.3) yield for all $t\in[0,T]$:
@@
We use the method of continuity [undefu, Lem. 16.2.2]. Define $A_{0}\in\mathcal{L}(V,V^{*})$ by $A_{0}v\coloneqq\theta(\cdot,v)_{V}$. For $\lambda\in[0,1]$, put
@@
Clearly, $L_{\lambda}$ is linear. We show that $L_{\lambda}\in\mathcal{L}(E,S\times H)$ and that $0,1]\to\mathcal{L}(E,S\times H)\colon\lambda\mapsto L_{\lambda}$ is continuous.
@@
$\|A_{0}(\cdot,w(\cdot))u(\cdot)\|_{L^{2}(0,T;V^{*})}\leq C_{n,T}\|u\|_{L^{2}(0,T;V)}<\infty$
@@
where $\alpha\coloneqq C_{n,T}(1+\|\psi\|_{L^{2}(0,T;U))})$.
@@
As $u|_{[0,\tilde{T}]}$ is a strong solution to (3.1) on $0,\tilde{T}$], (3.4) and (3.9) yield